# Towards Universal Soccer Video Understanding

### v2

@@ Appendix B SoccerNet-pro Dataset Details
As discussed in Section 3.2, alongside the SoccerReplay-1988 dataset, we incorporate two existing datasets, SoccerNet-v2 [9] and MatchTime [40], into the training process to enrich the training data. These datasets undergo the following preprocessing strategies and are then unified into the SoccerNet-pro dataset, ensuring consistency in format with SoccerReplay-1988.
@@ B.1 SoccerNet-v2
Table 8: Processing strategy for Mapping SoccerNet-pro Labels. The Reference column describes the specific processing applied to the original 17 classes of labels.
@@ B.2 MatchTime
The MatchTime dataset [40], curated from the SoccerNet-Caption [34], contains a substantial amount of commentary, with only a small portion accompanied by event labels. To bridge this gap, we apply the prompt-based approach described in Section A.3 to summarize the commentaries into event labels, assigning each commentary a corresponding label.
@@ B.3 Data Split Strategy
As described in Section 3.1, the SoccerReplay-1988 dataset is divided into 1,488 matches for training, 250 for validation, and 250 for testing. For the processed SoccerNet-pro dataset (including SoccerNet-v2 and MatchTime), we adhere to the original partitioning strategies and match distributions of its source datasets as detailed in Table 9.
@@ Appendix C Implementation Details
In this section, we provide additional implementation details about MatchVision. Section C.1 presents more information on data preprocessing strategies; Section C.2 elaborates on the evaluation strategies used during model training; and Section C.3 discusses several hyperparameter choices inspired by prior works.
@@ C.1 Data Preprocessing
Our automated data curation pipeline filters out video clips with missing annotations, incorrect cropping, or invalid timestamps. In all experiments, including pretraining and downstream tasks, video frames are resized to $224\times 224$ and preprocessed using the SigLIP [57] image preprocessor, which normalizes frames to a mean of 0.5 and a standard deviation of 0.5 before serving as the inputs. For overlapping video content between the SoccerNet-v2 [9] and MatchTime [40] datasets, we prioritize using event labels from SoccerNet-v2 for training.
@@ C.2 Validation Strategy during Training
For all experiments, we select the best-performing checkpoints on the validation set using appropriate evaluation strategies, as outlined below.
During MatchVision pretraining: (i) For supervised training, we adopt the top-1/3/5 event classification accuracy on the validation set to select the best model; (ii) For visual-language contrastive learning, video-to-text retrieval is performed, and the top-1/3/5 accuracy of event classification, comparing the retrieved text’s event labels to the ground truth, serves as the validation metric.
@@ C.3 Hyperparameter Selection
Temporal Window Size. We adopt a 30-second temporal window to extract video clips. This choice is inspired by MatchTime [40], which demonstrates that a 30-second window is sufficient to capture adequate visual information for optimal performance, outperforming the 45-second window used in SoccerNet-Caption [34].
LoRA Rank. For finetuning the commentary generation head, we use LoRA [23] with a rank of 16, following the approach outlined in MatchTime [40].
@@ Appendix D Experiments
In this section, we provide additional details to offer deeper insights into our model and its performance. Specifically, Section D.1 presents training curves to clearly illustrate the training process; and Section D.2 showcases more qualitative results, demonstrating the model’s capability to effectively understand soccer dynamics.
@@ D.1 Training Curves
We present the loss curves for the visual encoder pretraining in Figure 6. Our MatchVision demonstrates significantly better convergence compared to the SigLIP backbone, indicating that it effectively leverages spatiotemporal attention to utilize temporal information, learning representations better suited for highly dynamic soccer videos.
D.2 More Qualitative Results
In this part, we provide more qualitative visualizations of commentary generation across various events on the field, as depicted in Figure 78, and 9.
@@ Appendix E Limitations & Future Work
Although MatchVision explores to establish a universal visual-language foundation model tailored for soccer, it is not without its limitations: (i) Currently, MatchVision is adapted to event classification, commentary generation, and foul recognition tasks. In the future, we plan to further extend it to tasks such as player tracking and soccer game state reconstruction, aiming to develop a more comprehensive foundation model for soccer analysis. (ii) Following prior works, our commentary generation remains anonymized. This is left for future work, where we aim to fully leverage the contextual information available in our SoccerReplay-1988 dataset to enable more vivid, accurate, and context-aware commentary generation.

### v3

@@ Appendix B SoccerNet-pro Dataset Details
As discussed in the main text, alongside the SoccerReplay-1988 dataset, we also incorporate two existing datasets, SoccerNet-v2 [10] and MatchTime [41] to enrich the training data. These datasets undergo the following preprocessing strategies and are then unified into the SoccerNet-pro dataset, ensuring format consistency with SoccerReplay-1988.
@@ B.1 SoccerNet-v2
Table 8: Processing Strategy for SoccerNet-pro. The Reference column details the specific processing applied to the original labels.
@@ B.2 MatchTime
The MatchTime dataset [41], curated from SoccerNet-Caption [35], contains a substantial amount of commentary, with only a small portion accompanied by event labels. To bridge this gap, we apply the prompt-based approach described in Sec. A.3 to summarize commentaries into event labels, assigning each commentary a corresponding label.
@@ B.3 Data Split Strategy
As described in the manuscript, SoccerReplay-1988 is divided into 1,488 matches for training, 250 for validation, and 250 for testing. For the processed SoccerNet-pro dataset (including SoccerNet-v2 and MatchTime), we adhere to the original partitioning strategies and match distributions of its source datasets as detailed in Table 9.
@@ Appendix C Implementation Details
In this section, we provide additional implementation details about MatchVision. Sec. C.1 presents more information on data preprocessing strategies; Sec. C.2 elaborates on the evaluation strategies used during model training; and Sec. C.3 discusses several hyperparameter choices inspired by prior works.
@@ C.1 Data Preprocessing
Our automated data curation pipeline filters out video clips with missing annotations, incorrect cropping, or invalid timestamps. In all experiments, video frames are resized to $224\times 224$ and preprocessed using the image preprocessor of SigLIP [60], which normalizes frames to a mean of 0.5 and a standard deviation of 0.5 before serving as inputs. For overlapping video content between SoccerNet-v2 [10] and MatchTime [41], we prioritize using event labels from SoccerNet-v2.
@@ C.2 Validation Strategy during Training
We select the best-performing checkpoints on the validation set with the evaluation strategies detailed below:
During pretraining: (i) For supervised classification, we adopt top-1/3/5 event classification accuracy on the validation set to select the best model; (ii) For visual-language contrastive learning, video-to-text retrieval is performed, with top-1/3/5 accuracy of event classification (comparing retrieved texts’ event labels to ground truth) as the validation metric.
@@ C.3 Hyperparameter Selection
Temporal Window Size. We adopt a 30-second temporal window to extract video clips. This is inspired by MatchTime [41], which demonstrates that a 30-second window is sufficient to capture adequate visual information for optimal performance, outperforming the 45-second window used in SoccerNet-Caption [35].
LoRA Rank. For finetuning the commentary generation head, we use LoRA [24] with a rank of 16, following [41].
@@ Appendix D Experiments
In this section, we provide additional details to offer deeper insights into our model and its performance. Specifically, Sec. D.1 presents training curves to clearly illustrate the training process; Sec. D.2 and Sec. D.3 showcase more quantitative and qualitative results, respectively, demonstrating the model’s capability to effectively understand soccer dynamics.
@@ D.1 Training Curves
We present the loss curves for visual encoder pretraining in Figure 6. Our MatchVision demonstrates significantly better convergence compared to SigLIP [60] backbone, indicating that it effectively leverages spatiotemporal attention to utilize temporal information, learning representations better suited for highly dynamic soccer videos.
D.2 More Quantitative Results
Here, we compare present comparisons with more advanced methods in the SoccerNet Foul Recognition challenge [9], where MatchVision remains competitive even with a frozen visual encoder.
Top-1 Accuracy Ours Baseline zyz PD_PS_GSN Redsox xiao_he_shang
Foul 0.44 0.36 0.58 0.44 0.60 0.46
Severity 0.58 0.54 0.58 0.50 0.05 0.47
Table 10: More Quantitative Results on Multi-view Foul Recognition.
D.3 More Qualitative Results
More qualitative visualizations of commentary generation across various events on the field are depicted in Figure 78, and 9.
@@ Appendix E Limitations & Future Work
Although MatchVision explores establishing a soccer-specific visual encoder, it is not without its limitations: (i) Currently, MatchVision is adapted to event classification, commentary generation, and foul recognition tasks. In the future, we plan to further extend it to more challenging tasks such as player tracking and dense video captioning, aiming to develop a more comprehensive foundation model for soccer analysis. (ii) Given computational and annotation constraints, SoccerReplay-1988 primarily focuses on European league soccer data. We aim to leverage our scalable automated annotation pipeline to further expand the dataset, encompassing a more comprehensive range of soccer data. (iii) Following prior works, our commentary generation remains anonymized. This is left for future work, where we aim to fully leverage contextual information available in our SoccerReplay-1988 dataset to enable more vivid, accurate, and context-aware commentary generation.